# The Bisognano-Wichmann property for non-unitary Wightman conformal field theories

### v1

@@ 5 Haag duality for Wightman CFTs with invariant Hermitian forms
In this section we continue the notation of Section 4, fixing an involutive Wightman CFT. We will now study the notion of ‘Haag duality’ for nets of algebras, which is a key application of Tomita-Takesaki theory in the unitary setting. Our first step will be to make contact with the theory of standard subspaces, which have proven to be very powerful in analyzing duality phenomena in algebraic quantum field theory [Lon08a, Lon08b] (see also Section 2.1 for the unitary case). The results in this section on Haag duality for non-unitary theories are exploratory at this stage, with the main result being some sufficient conditions for Haag duality of the double commutant net ${\mathcal{P}}(I)^{\prime\prime}$, albeit conditions that may not be practical to check. However, given how crucial Haag duality is in the study of nets of operator algebras, it is worthwhile to initiate the study in this non-unitary setting. Furthermore, the results of this section will be applied in Section 6 to gain insight into unitary models.
@@
Let ${\mathcal{D}}$ be a locally convex space with nondegenerate Hermitian form, equipped with one of its standard topologies. Then the correspondence $K\leftrightarrow S_{K}$ induces a bijection between i) standard subspaces $K$ of ${\mathcal{D}}$, and 2) closed conjugate-linear involutions $S$
@@
This is the same as the perpendicular subspace to $iK$ with respect to the real form $\operatorname{Re}\langle\,\cdot\,,\,\cdot\,\rangle$.
@@
Hence if $\Psi\in K^{\prime}$ we see $\langle S_{K}\Phi,\Psi\rangle=\langle\Psi,\Phi\rangle$, which shows that $\Psi\in D(S_{K})^{*}$ and $S_{K}^{*}\Psi=1$, i.e., $\Psi\in\tilde{K}$. Conversly, if $\Psi\in\tilde{K}$ then $\langle S_{K}\Phi,\Psi\rangle=\langle\Psi,\Phi\rangle$, and we see that $\langle\Psi,\Phi\rangle\in{\mathbb{R}}$, so that $\Psi\in K^{\prime}$. We conclude $K^{\prime}=\tilde{K}$, completing the proof of (ii). ∎
@@
Let ${\mathcal{D}}$ be a locally convex space equipped with a jointly continuous nondegenerate Hermitian form and a strongly continuous untiary representation $U:\operatorname{M\ddot{o}b}\to\operatorname{End}({\mathcal{D}})$. A (non-unitary) Möbius-covariant net of standard subspaces is a collection of standard subspaces ${\mathcal{K}}(I)\subset{\mathcal{D}}$ such that the following holds:
@@
In the context of positive energy nets of standard subspaces of a Hilbert space, Haag duality is automatic [Lon08a, Thm. 3.3.1]. We now investigate the question of Haag duality for the nets of standard subspaces arising from a Wightman CFT.
@@
    If ${\mathcal{L}}_{\pm}$ is the standard subspace corresponding to $S_{\pm}$, then ${\mathcal{L}}_{\pm}^{\prime}={\mathcal{L}}_{\mp}$, ${\mathcal{K}}(I_{\pm})\subset{\mathcal{K}}_{w}(I_{\pm})\subset{\mathcal{L}}_{\pm}$, and $V_{t}{\mathcal{L}}_{\pm}={\mathcal{L}}_{\pm}$ for all $t\in{\mathbb{R}}$.
@@
We first show that $\theta$ maps $D(V_{i\pi})$ into $D(V_{-i\pi})$. Indeed if $\Phi\in D(V_{i\pi})$ and $\Phi^{\prime}\in{\mathcal{D}}$, then we have a map $G_{\Phi^{\prime}}:{\mathbb{S}}_{\pi}\to{\mathbb{C}}$ that is continuous and holomorphic in the interior, and
$G_{\Phi^{\prime}}(t)=\langle V_{t}\Phi,\Phi^{\prime}\rangle,\qquad G_{\Phi^{\prime}}(t+i\pi)=\langle V_{t}V_{i\pi}\Phi,\Phi^{\prime}\rangle.$
Define $H_{\Phi^{\prime}}:{\mathbb{S}}_{-\pi}\to{\mathbb{C}}$ by $H_{\Phi^{\prime}}(\tau)=\overline{G_{\theta\Phi^{\prime}}(\overline{\tau})}$. Since $\theta$ commutes with $V_{t}$, we have
@@
Hence $\theta\Phi\in D(V_{-i\pi})$ and $V_{-i\pi}\theta\Phi=\theta V_{i\pi}\Phi$. The same argument shows that if $\Phi\in D(V_{-i\pi})$ then $\theta\Phi\in D(V_{i\pi})$ and $\theta V_{-i\pi}\Phi=V_{i\pi}\theta\Phi$. Since $\theta$ is an involution and maps $D(V_{\pm i\pi})$ into $D(V_{\mp i\pi})$, we actually must have $\theta D(V_{\pm i\pi})=D(V_{\mp i\pi})$, and so from our previous calculations $\theta V_{\pm i\pi}\theta=V_{\mp i\pi}$. A very similar calculation, this time defining $H_{\Phi^{\prime}}(\tau)=G_{\Phi^{\prime}}(\tau-i\pi)$, shows that $V_{\pm i\pi}V_{\mp i\pi}=\mathrm{Id}_{D(V_{\mp i\pi})}$, completing (i).
From part (i), we see that $S_{\pm}^{2}=V_{\mp i\pi}V_{\pm i\pi}=\mathrm{Id}_{D(\pm i\pi)}=\mathrm{Id}_{D(S_{\pm})}$, and so $S_{\pm}$ are conjugate-linear involutions. The maps $V_{\pm i\pi}$ are closed for the ${\mathcal{D}}$-weak topology (and so for every standard topology) since adjoints are automatically ${\mathcal{D}}$-weakly closed and $V_{\pm i\pi}$ are self-adjoint. Hence $S_{\pm}$ are closed as well, and we have
@@
The identity ${\mathcal{L}}_{\pm}^{\prime}={\mathcal{L}}_{\mp}$ now follows from Lemma 5.5. By definition ${\mathcal{L}}_{\pm}$ consists of fixed points of $S_{\pm}$, and so ${\mathcal{P}}(I_{\pm})_{sa}\Omega\subset{\mathcal{L}}_{\pm}$ by Theorem 4.2. Taking closures yields ${\mathcal{K}}(I_{\pm})\subset{\mathcal{L}}_{\pm}$. Finally note that $V_{t}$ commutes with both $\theta$ and $V_{\pm i\pi}$, and so with $S_{\pm}$. Hence $V_{t}$ leaves ${\mathcal{L}}_{\pm}$ invariant for all $t$, which implies that $V_{t}{\mathcal{L}}_{\pm}={\mathcal{L}}_{\pm}$. ∎
@@
If ${\mathcal{P}}(I_{+})\Omega$ is a ${\mathcal{F}}$-strong core for $V_{i\pi}$ then ${\mathcal{K}}(I_{+})={\mathcal{L}}_{+}$, and applying $\theta$ yields ${\mathcal{K}}(I_{-})={\mathcal{L}}_{-}$. Hence by Lemma 5.12 we have ${\mathcal{K}}(I_{\pm})^{\prime}={\mathcal{K}}(I_{\mp})$. Haag duality then follows from Möbius covariance. For the converse, suppose the net is Haag dual, which is to say ${\mathcal{K}}(I_{\pm})^{\prime}={\mathcal{K}}(I_{\mp})$. By Lemma 5.12 we have
${\mathcal{K}}(I_{\pm})\subset{\mathcal{L}}_{\pm}={\mathcal{L}}_{\mp}^{\prime}\subset{\mathcal{K}}(I_{\mp})^{\prime}={\mathcal{K}}(I_{\pm}).$
As the first and last terms coincide, we have ${\mathcal{K}}(I_{\pm})={\mathcal{L}}_{\pm}$ and so ${\mathcal{P}}(I_{+})$ is a ${\mathcal{F}}$-strong core for $V_{i\pi}$. The proof for the ${\mathcal{D}}$-weak topology is similar. ∎
Remark 5.14.
We see from the proof of Corollary 5.13 that the key question for (weak) Haag duality is whether the inclusion ${\mathcal{K}}(I_{\pm})\subset{\mathcal{L}}_{\pm}$ is an equality. This is an inclusion of standard subspaces, the involution $S_{\pm}$ satisfies $S_{\pm}=\theta V_{\pm i\pi}$, and ${\mathcal{K}}(I_{\pm})$ is invariant under $V_{t}$. In the unitary setting the identity ${\mathcal{K}}(I_{\pm})={\mathcal{L}}_{\pm}$ would follow (see e.g. [Lon08a, Prop. 2.1.10]), but we do not see an immediate generalization to our context.
Remark 5.15.
In light of Lemma 5.12, one could alternatively try to construct a Haag dual Möbius covariant net of standard subspaces as follows. Given an interval $I$, choose $\gamma\in\operatorname{M\ddot{o}b}$ such that $\gamma(I_{+})=I$, and set ${\mathcal{L}}(I)=U(\gamma){\mathcal{L}}_{+}$. Since $\gamma(I_{+})=I_{+}$ exactly when $\gamma=v_{t}$ for some $t$, and $V_{t}{\mathcal{L}}_{+}={\mathcal{L}}_{+}$, the subspace ${\mathcal{L}}(I)$ does not depend on the choice of $\gamma$, and in particular ${\mathcal{L}}(I_{+})={\mathcal{L}}_{+}$. If $r_{\pi}\in\operatorname{M\ddot{o}b}$ is rotation by $\pi$, then $U(r_{\pi})V_{t}U(r_{\pi})^{-1}=V_{-t}$ and it is then straightforward to check that $U(r_{\pi})V_{i\pi}U(r_{\pi})^{-1}=V_{-i\pi}$. Hence ${\mathcal{L}}(I_{-})={\mathcal{L}}_{-}$ as well. This collection of standard subspaces is Haag dual (in particular, local) and Möbius covariant, but does not visibly satisfy isotony.
We conclude this section with a discussion of Haag duality for nets of algebras. As with standard subspaces, we have a preliminary definition of a local net of algebras (for which one does not expect a robust general theory, in part because of the lack of a positive energy condition).
@@
Let ${\mathcal{D}}$ be a locally convex space equipped with a jointly continuous nondegenerate Hermitian form, a strongly continuous untiary representation $U:\operatorname{M\ddot{o}b}\to\operatorname{End}({\mathcal{D}})$, and a choice of non-zero vector $\Omega\in{\mathcal{D}}$. A (non-unitary) Möbius-covariant net of algebras is a collection of unital *-subalgebras ${\mathcal{Q}}(I)\subset\operatorname{End}_{*}({\mathcal{D}})$ such that the following holds:
@@
We will verify that Haag duality occurs under a pair of reasonable-looking (but presumably difficult to verify) conditions. The two arguments share an important technical (but standard) step, which we distil into the following lemma.
@@
We can now give sufficient conditions for Haag duality of the net ${\mathcal{Q}}(I)$:
Proposition 5.18.
Let ${\mathcal{Q}}(I)$ be the Möbius-covariant net of algebras ${\mathcal{Q}}(I)={\mathcal{P}}(I)^{\prime\prime}$. Then any one of the following conditions individually implies that ${\mathcal{Q}}$ is Haag dual:
  1. (a)
    The net of standard subspaces ${\mathcal{K}}(I)$ is weakly Haag dual, i.e. ${\mathcal{K}}_{w}(I^{\prime})={\mathcal{K}}_{w}(I)^{\prime}$
  2. (b)
    ${\mathcal{P}}(I_{+})\Omega$ is a core for $V_{i\pi}$ for any standard topology
  3. (c)
    ${\mathcal{P}}(I_{+})^{\prime\prime}=\overline{{\mathcal{P}}(I_{+})}$ where the closure is taken in the ${\mathcal{D}}$-weak operator topology [^10] That is, the topology of ${\mathcal{D}}$-weak pointwise convergence..
@@
First consider assumption (a). We know that $\theta{\mathcal{P}}(I_{+})\theta={\mathcal{P}}(I_{-})$, and taking double commutants yields $\theta{\mathcal{Q}}(I_{+})\theta={\mathcal{Q}}(I_{-})$. Thus to establish Haag duality, we must show $\theta{\mathcal{Q}}(I_{+})\theta={\mathcal{Q}}(I_{+})^{\prime}$. The locality axiom implies that $\theta{\mathcal{Q}}(I_{+})\theta\subset{\mathcal{Q}}(I_{+})^{\prime}$, so we need only show the reverse inclusion. Plugging in ${\mathcal{Q}}(I)={\mathcal{P}}(I)^{\prime\prime}$, we wish to show ${\mathcal{P}}(I_{+})^{\prime\prime\prime}\subset\theta{\mathcal{P}}(I_{+})^{\prime\prime}\theta$. It suffices to show that $\theta{\mathcal{P}}(I_{+})^{\prime}\theta\subset{\mathcal{P}}(I_{+})^{\prime\prime}$, and then take commutants.
We will establish this last inclusion by invoking Lemma 5.17 with $A={\mathcal{P}}(I_{+})^{\prime}$ and $B={\mathcal{P}}(I_{+})$. The conclusion of the lemma is exactly the required inclusion, so we just verify that the two hypotheses hold. We first consider hypothesis (i) of the Lemma. Note that if $X\in{\mathcal{P}}(I_{+})^{\prime}$ and $X=X^{*}$, then for all $Y\in{\mathcal{P}}(I_{+})_{sa}$ we have $\langle X\Omega,Y\Omega\rangle\in{\mathbb{R}}$, and so $X\Omega\in{\mathcal{K}}(I_{+})^{\prime}={\mathcal{K}}_{w}(I_{-})$ by assumption (a). Similarly $({\mathcal{P}}(I_{-})^{\prime})_{sa}\Omega\subset{\mathcal{K}}_{w}(I_{+})$, and so $({\mathcal{P}}(I_{+})^{\prime})_{sa}\Omega$ and $({\mathcal{P}}(I_{-})^{\prime})_{sa}\Omega$ are in particular local. If $X,Y\in({\mathcal{P}}(I_{+})^{\prime})_{sa}$, then $\theta X\theta\in{\mathcal{P}}(I_{-})^{\prime}$ so

### v2

@@ 5 Haag duality for Wightman CFTs with invariant Hermitian forms
In this section we continue the notation of Section 4, fixing an involutive Wightman CFT. We will now study the notion of ‘Haag duality’ for nets of algebras, which is a key application of Tomita-Takesaki theory in the unitary setting. Our first step will be to make contact with the theory of standard subspaces, which have proven to be very powerful in analyzing duality phenomena in algebraic quantum field theory [Lon08a, Lon08b] (see also Section 2.1 for the unitary case). The main result of this section is the Haag duality of the double commutant net ${\mathcal{P}}(I)^{\prime\prime}$. The results of this section will also be applied in Section 6 to gain insight into unitary models.
@@
Let $\mathcal{D}$} be a locally convex space with nondegenerate Hermitian form, equipped with one of its standard topologies. Then the correspondence $K\leftrightarrow S_{K}$ induces a bijection between i) standard subspaces $K$ of $\mathcal{D}$}, and ii) densely-defined closed conjugate-linear involutions $S$.
@@
This is the same as the perpendicular subspace to $iK$ with respect to the real form $\operatorname{Re}\langle\,\cdot\,$$\cdot\,\rangle$.
@@
Hence if $\Psi\in K^{\prime}$ we see $\langle S_{K}\Phi,\Psi\rangle=\langle\Psi,\Phi\rangle$, which shows that $\Psi\in D(S_{K}^{*})$ and $S_{K}^{*}\Psi=\Psi$, i.e., $\Psi\in\tilde{K}$. Conversely, if $\Psi\in\tilde{K}$ then $\langle S_{K}\Phi,\Psi\rangle=\langle\Psi,\Phi\rangle$, and we see that $\langle\Psi,\Phi\rangle\in{\mathbb{R}}$, so that $\Psi\in K^{\prime}$. We conclude $K^{\prime}=\tilde{K}$, completing the proof of (ii). ∎
@@
Let $\mathcal{D}$} be a locally convex space equipped with a jointly continuous nondegenerate Hermitian form and a strongly continuous unitary representation $U:\operatorname{M\ddot{o}b}\to\operatorname{End}({\mathcal{D}})$. A (non-unitary) Möbius-covariant net of standard subspaces is a collection of standard subspaces ${\mathcal{K}}(I)\subset{\mathcal{D}}$ such that the following holds:
@@
In the context of positive energy nets of standard subspaces of a Hilbert space, Haag duality is automatic [Lon08a, Thm. 3.3.1]. We now investigate the question of Haag duality for the nets of standard subspaces ${\mathcal{K}}(I)$ and ${\mathcal{K}}_{w}(I)$ arising from a Wightman CFT (as in Definition 5.6).
@@
    If $\mathcal{L}$$_{\pm}$ is the standard subspace corresponding to $S_{\pm}$, then $\mathcal{L}$$_{\pm}^{\prime}$$\mathcal{L}$$_{\mp}$, ${\mathcal{K}}(I_{\pm})\subset{\mathcal{K}}_{w}(I_{\pm})\subset{\mathcal{L}}_{\pm}$, and $V_{t}{\mathcal{L}}_{\pm}={\mathcal{L}}_{\pm}$ for all $t\in{\mathbb{R}}$.
  4. (iv)
    ${\mathcal{K}}(I_{\pm})^{\prime}={\mathcal{K}}_{w}(I_{\pm})^{\prime}={\mathcal{L}}_{\mp}$
@@
We first show that $\theta$ maps $D(V_{i\pi})$ into $D(V_{-i\pi})$. Indeed if $\Phi\in D(V_{i\pi})$ and $\Phi^{\prime}\in{\mathcal{D}}$, then we have a map $G_{\Phi^{\prime}}:{\mathbb{S}}_{i\pi}\to{\mathbb{C}}$ that is continuous and holomorphic in the interior, and
$G_{\Phi^{\prime}}(t)=\langle V_{t}\Phi,\Phi^{\prime}\rangle,\qquad G_{\Phi^{\prime}}(t+i\pi)=\langle V_{t}V_{i\pi}\Phi,\Phi^{\prime}\rangle.$ (5.1)
Define $H_{\Phi^{\prime}}:{\mathbb{S}}_{-i\pi}\to{\mathbb{C}}$ by $H_{\Phi^{\prime}}(\tau)=\overline{G_{\theta\Phi^{\prime}}(\overline{\tau})}$. Since $\theta$ commutes with $V_{t}$, we have
@@
Hence $\theta\Phi\in D(V_{-i\pi})$ and $V_{-i\pi}\theta\Phi=\theta V_{i\pi}\Phi$. The same argument shows that if $\Phi\in D(V_{-i\pi})$ then $\theta\Phi\in D(V_{i\pi})$ and $\theta V_{-i\pi}\Phi=V_{i\pi}\theta\Phi$. Since $\theta$ is an involution and maps $D(V_{\pm i\pi})$ into $D(V_{\mp i\pi})$, we actually must have $\theta D(V_{\pm i\pi})=D(V_{\mp i\pi})$, and so from our previous calculations $\theta V_{\pm i\pi}\theta=V_{\mp i\pi}$.
We next show that $V_{i\pi}^{-1}=V_{-i\pi}$, or equivalently that $V_{\mp i\pi}V_{\pm i\pi}=\mathrm{Id}_{D(V_{\pm i\pi})}$. We will show $V_{-i\pi}V_{i\pi}=\mathrm{Id}_{D(V_{i\pi})}$, and the other identity is similar. Let $\Phi\in D(V_{i\pi})$ and $\Phi^{\prime}\in{\mathcal{D}}$, and let $G_{\Phi^{\prime}}$ be as in (5.1). We now define $\widetilde{H}_{\Phi^{\prime}}:{\mathbb{S}}_{-i\pi}\to{\mathbb{C}}$ by $\widetilde{H}_{\Phi^{\prime}}(\tau)=G_{\Phi^{\prime}}(\tau+i\pi)$. Then for $t\in{\mathbb{R}}$ we have
$\widetilde{H}_{\Phi^{\prime}}(t)=\langle V_{t}V_{i\pi}\Phi,\Phi^{\prime}\rangle,\qquad\widetilde{H}_{\Phi^{\prime}}(t-i\pi)=\langle V_{t}\Phi,\Phi^{\prime}\rangle.$
Since $\Phi^{\prime}$ was arbitrary, this demonstrates that $V_{i\pi}\Phi\in D(V_{-i\pi})$ and that $V_{-i\pi}V_{i\pi}\Phi=\Phi$, as claimed. This completes the proof of (i).
From part (i), we see that $S_{\pm}^{2}=V_{\mp i\pi}V_{\pm i\pi}=\mathrm{Id}_{D(V_{\pm i\pi})}=\mathrm{Id}_{D(S_{\pm})}$, and so $S_{\pm}$ are conjugate-linear involutions. The maps $V_{\pm i\pi}$ are closed for the $\mathcal{D}$}-weak topology (and so for every standard topology) since adjoints are automatically $\mathcal{D}$}-weakly closed and $V_{\pm i\pi}$ are self-adjoint by Theorem 4.6. Hence $S_{\pm}$ are closed as well, and we have
@@
The identity $\mathcal{L}$$_{\pm}^{\prime}$$\mathcal{L}$$_{\mp}$ now follows from Lemma 5.5. By definition $\mathcal{L}$$_{\pm}$ consists of fixed points of $S_{\pm}$, and so ${\mathcal{P}}(I_{\pm})_{sa}\Omega\subset{\mathcal{L}}_{\pm}$ by Theorem 4.2. Taking closures yields ${\mathcal{K}}(I_{\pm})\subset{\mathcal{L}}_{\pm}$. Finally note that $V_{t}$ commutes with both $\theta$ and $V_{\pm i\pi}$, and so with $S_{\pm}$. Hence $V_{t}$ leaves $\mathcal{L}$$_{\pm}$ invariant for all $t$, which implies that $V_{t}{\mathcal{L}}_{\pm}={\mathcal{L}}_{\pm}$, completing the proof of (iii).
To prove (iv), it suffices to show that ${\mathcal{K}}(I_{\pm})^{\prime}={\mathcal{L}}_{\mp}$, and by (iii) we only need to check ${\mathcal{K}}(I_{\pm})^{\prime}\subseteq{\mathcal{L}}_{\mp}$. Without loss of generality we only consider the inclusion ${\mathcal{K}}(I_{+})^{\prime}\subseteq{\mathcal{L}}_{-}$. Let $\Phi\in{\mathcal{K}}(I_{+})^{\prime}$. Then for all $\Psi\in{\mathcal{K}}(I_{+})$ we have
$\langle\Psi,\Phi\rangle=\langle\Phi,\Psi\rangle=\langle\Phi,S_{+}\Psi\rangle=\langle\Phi,\theta V_{i\pi}\Psi\rangle=\langle V_{i\pi}\Psi,\theta\Phi\rangle$
where the first equality is the definition of ${\mathcal{K}}(I_{+})^{\prime}$, the second is the fact that ${\mathcal{K}}(I_{+})\subset{\mathcal{L}}_{+}$, the third is the definition of $S_{+}$, and the fourth is the fact that $\theta$ is an antiunitary involution. The identity
$\langle\Psi,\Phi\rangle=\langle V_{i\pi}\Psi,\theta\Phi\rangle$ (5.2)
extends by linearity to all $\Psi\in{\mathcal{K}}(I_{+})+i{\mathcal{K}}(I_{+})$, and in particular (5.2) holds for all $\Psi\in{\mathcal{P}}(I_{+})\Omega$. Hence $\theta\Phi$ lies in the domain of $(V_{i\pi}|_{{\mathcal{P}}(I_{+})\Omega})^{*}$, and we recall from Theorem 4.6 that $(V_{i\pi}|_{{\mathcal{P}}(I_{+})\Omega})^{*}=V_{i\pi}$. Thus from (5.2) we may deduce that $\theta\Phi\in D(V_{i\pi})$ and that $\langle\Psi,\Phi\rangle=\langle\Psi,V_{i\pi}\theta\Phi\rangle$ for all $\Psi\in{\mathcal{P}}(I_{+})\Omega$. We then have
$\Phi=V_{i\pi}\theta\Phi=\theta V_{-i\pi}\Phi=S_{-}\Phi.$
where the first equality uses the Reeh-Schlieder property and the nondegeneracy of the form, the second uses part (i) and implicitly asserts that $\Phi\in D(V_{-i\pi})$, and the third is the definition of $S_{-}$. Hence by definition we have $\Phi\in{\mathcal{L}}_{-}$, as required. ∎
As a consequence of Lemma 5.12, we can show that the ‘dual net’ $I\mapsto{\mathcal{K}}(I^{\prime})^{\prime}$ is local, and moreover satisfies Haag duality.
Corollary 5.13.
The net of subspaces ${\mathcal{L}}(I):={\mathcal{K}}(I^{\prime})^{\prime}$ is a Möbius-covariant net of standard subspaces of $\mathcal{D}$}, and moreover this net satisfies Haag duality.
Proof.
Isotony and Möbius-covariance of the net ${\mathcal{L}}(I)$ follow from the corresponding properties of the net ${\mathcal{K}}(I)$. By Möbius covariance, it suffices to check locality and Haag duality for the pair of intervals $I_{\pm}$. Observe that ${\mathcal{L}}(I_{\pm})$ is the space $\mathcal{L}$$_{\pm}$ of Lemma 5.12, by part (iv) of that lemma. We have $\mathcal{L}$$_{\pm}^{\prime}$$\mathcal{L}$$_{\mp}$ by part (iii) of the same lemma, which immediately implies locality and Haag duality. ∎
Remark 5.14.
We have ${\mathcal{K}}(I)\subseteq{\mathcal{L}}(I)$ by the locality of $\mathcal{K}$}, and evidently the Haag duality of the net $\mathcal{K}$} is equivalent to this inclusion being an equality. By Haag duality of the net $\mathcal{L}$}, we have ${\mathcal{L}}(I)={\mathcal{K}}(I)^{\prime\prime}$, so the question of Haag duality of the net $\mathcal{K}$} rests on the difference between taking the closure of a subspace versus taking the double complement. The question of Haag duality for $\mathcal{K}$} can also be interpreted in the context of inclusions of standard subspaces ${\mathcal{K}}(I)\subset{\mathcal{L}}(I)$. The involution $S_{\pm}$ corresponding to ${\mathcal{L}}(I_{\pm})$ satisfies $S_{\pm}=\theta V_{\pm i\pi}$, and ${\mathcal{K}}(I_{\pm})$ is invariant under $V_{t}$. In the unitary setting the identity ${\mathcal{K}}(I_{\pm})={\mathcal{L}}(I_{\pm})$ would follow (see e.g. [Lon08a, Prop. 2.1.10]), but we do not see an immediate generalization to our context.
In light of Lemma 5.12, it is also possible to rephrase the question of Haag duality for $\mathcal{K}$} in operator theoretic language.
@@
If ${\mathcal{P}}(I_{+})\Omega$ is a $\mathcal{F}$}-strong core for $V_{i\pi}$ then ${\mathcal{K}}(I_{+})={\mathcal{L}}(I_{+})$. By Möbius-covariance we have ${\mathcal{K}}(I)={\mathcal{L}}(I)$ for all intervals $I$, and Haag duality follows. For the converse, suppose the net $\mathcal{K}$} is Haag dual, in which case ${\mathcal{K}}(I_{+})={\mathcal{K}}(I_{-})^{\prime}={\mathcal{L}}(I_{+})$. It follows that ${\mathcal{P}}(I_{+})\Omega$ is a $\mathcal{F}$}-strong core for $V_{i\pi}$. The proof for the $\mathcal{D}$}-weak topology is similar. ∎
We conclude this section with a study of Haag duality for nets of algebras. As with standard subspaces, we have a preliminary definition of a local net of algebras (for which one does not expect a robust general theory, in part because of the lack of a positive energy condition).
@@
Let $\mathcal{D}$} be a locally convex space equipped with a jointly continuous nondegenerate Hermitian form, a strongly continuous unitary representation $U:\operatorname{M\ddot{o}b}\to\operatorname{End}({\mathcal{D}})$, and a choice of non-zero vector $\Omega\in{\mathcal{D}}$. A (non-unitary) Möbius-covariant net of algebras is a collection of unital *-subalgebras ${\mathcal{Q}}(I)\subset\operatorname{End}_{*}({\mathcal{D}})$ such that the following holds:
@@
We will shortly verify the Haag duality property for the net of algebras ${\mathcal{Q}}(I)$. First, however, we isolate an important technical step. The following lemma is analogous to standard computations in the context of Tomita-Takesaki theory and Haag duality for unitary theories.
@@
We can now prove Haag duality of the net ${\mathcal{Q}}(I)$:
Theorem 5.18 (Haag duality).
Let $({\mathcal{F}},{\mathcal{D}},U,\Omega)$ be an involutive Wightman CFT on $S^{1}$, and let ${\mathcal{P}}(I)\subset\operatorname{End}_{*}({\mathcal{D}})$ be the unital algebra generated by operators $\varphi(f)$, where $\varphi\in{\mathcal{F}}$ and $\operatorname{supp}f\subset I$. Let ${\mathcal{Q}}(I)$ be the Möbius-covariant net of algebras ${\mathcal{Q}}(I):={\mathcal{P}}(I)^{\prime\prime}$. Then $\mathcal{Q}$} satisfies the Haag duality condition: ${\mathcal{Q}}(I)^{\prime}={\mathcal{Q}}(I^{\prime})$ for all intervals $I$, where the commutant is taken in $\operatorname{End}_{*}({\mathcal{D}})$. Moreover, ${\mathcal{Q}}(I)={\mathcal{P}}(I^{\prime})^{\prime}$.
@@
By Möbius covariance, it suffices to consider the interval $I=I_{+}$. Observe that is suffices to prove the “moreover” statement ${\mathcal{Q}}(I_{+})={\mathcal{P}}(I_{-})^{\prime}$, as taking commutants in this identity immediately yields the Haag duality of $\mathcal{Q}$}. By locality of the net $\mathcal{P}$}, we have ${\mathcal{P}}(I_{-})\subseteq{\mathcal{P}}(I_{+})^{\prime}$, and taking commutants yields ${\mathcal{Q}}(I_{+})={\mathcal{P}}(I_{+})^{\prime\prime}\subseteq{\mathcal{P}}(I_{-})^{\prime}$. We therefore need only prove the reverse inclusion ${\mathcal{P}}(I_{-})^{\prime}\subseteq{\mathcal{P}}(I_{+})^{\prime\prime}$. We know that ${\mathcal{P}}(I_{-})=\theta{\mathcal{P}}(I_{+})\theta$, and taking commutants yields ${\mathcal{P}}(I_{-})^{\prime}=\theta{\mathcal{P}}(I_{+})^{\prime}\theta$. Thus the theorem reduces to establishing that $\theta{\mathcal{P}}(I_{+})^{\prime}\theta\subset{\mathcal{P}}(I_{+})^{\prime\prime}$.
We will establish this last inclusion by invoking Lemma 5.17 with $A={\mathcal{P}}(I_{+})^{\prime}$ and $B={\mathcal{P}}(I_{+})$. The conclusion of the lemma is exactly the required inclusion, so we just verify that the two hypotheses hold. We first consider hypothesis (i) of the Lemma. Note that if $X\in{\mathcal{P}}(I_{+})^{\prime}$ and $X=X^{*}$, then for all $Y\in{\mathcal{P}}(I_{+})_{sa}$ we have $\langle X\Omega,Y\Omega\rangle\in{\mathbb{R}}$, and so $X\Omega\in{\mathcal{K}}(I_{+})^{\prime}$.
If we set ${\mathcal{L}}(I_{\pm})={\mathcal{K}}(I_{\mp})^{\prime}$, then we have just shown that $({\mathcal{P}}(I_{+})^{\prime})_{sa}\Omega\subset{\mathcal{L}}(I_{-})$. A similar calculation shows that $({\mathcal{P}}(I_{-})^{\prime})_{sa}\Omega\subset{\mathcal{L}}(I_{+})$. By Corollary 5.13, the net ${\mathcal{L}}(I)$ is local, and so in particular the pair of subspaces $({\mathcal{P}}(I_{+})^{\prime})_{sa}\Omega$ and $({\mathcal{P}}(I_{-})^{\prime})_{sa}\Omega$ are local as well. If $X,Y\in({\mathcal{P}}(I_{+})^{\prime})_{sa}$, then $\theta X\theta\in{\mathcal{P}}(I_{-})^{\prime}$ and it now follows that